I need to solve for x.
4x(x-2)-5x(x-1)=2
I have already taken the steps to make the problem
4x^2-8x-5x^2-5x=2, but that is all I have. I have been trying by "substitution", but have not found the answer. Please help me solve for x.
Thank you,
Rachel

Answer by checkley77(12569)

4x(x-2)-5x(x-1)=2
4X^2-8X-5X^2+5X-2=0
-X^2-3X-2=0
X^2+3X+2=0
(X+2)(X+1)=0
X+2=0
X=-2 ANS.
X+1=0
X=-1 ANS.
PROOFS:
4*-2(-2-2)-5*-2(-2-1)=2
-8*-4+10*-3=2
32-30=2
2=2
4*-1(-1-2)-5*-1(-1-1)=2
-4*-3+5*-2=2
12-10=2